# Discovering Invariant Neighborhood Patterns for Non-Homophilous Graphs

### v1

@@ 2 Related work
Debiased Graph Neural Network. Debiased Graph Neural Networks aim to address the bias issue on graphs. To overcome degree-related bias, SL-DSGCN Tang et al. (2020) mitigates the degree-related bias of GCNs by capturing both discrepancies and similarities of nodes with different degrees. ImGAGN Qu et al. (2021) is proposed to address the class-related bias, ImGAGN generates a set of synthetic minority nodes to balance the class distribution. Different from these works on a single bias, BA-GNN Chen et al. (2022) proposes a novel Bias-Aware Graph Neural Network (BA-GNN) framework by learning node representation that is invariant across different biases and distributions for invariant prediction. EERM Wu et al. (2022) facilitates graph neural networks to leverage invariance principles for prediction, EERM resorts to multiple context explorers that are adversarially trained to maximize the variance of risks from multiple virtual environments, which enables the model to extrapolate from a single observed environment which is the common case for node-level prediction. However, these works heavily rely on the assumption of homophily, which may be prohibitively failed on non-homophilous graphs. Our method differs from the above methods and aims to learn invariant graph representation for non-homophilous graphs.
@@ 3 Discovering Invariant Neighborhood Patterns
Let $G=(V,E,X)$ be a graph, where $V$ is the set of nodes, $E$ is the set of edges, $X$in $R^{N\times d}$ is a feature matrix, each row in $X$ indicates a d-dimensional vector of a node. $A$in $\{0,1\}^{N\times N}$ is the adjacency matrix, where $A_{uv}=1$ if there exists an edge between node $v_{i}$ and $v_{j}$, otherwise $A_{ij}=0$ For semi-supervised node classification tasks, only part of nodes have known labels $Y^{o}=\{y_{1},y_{2},...y_{n}\}$, where $y_{j}\in\{0,1,...,c-1\}$ denotes the label of node $v_{j}$, $c$ is the number of classes. Similar to Chen et al. (2022), we also define a graph environment to be a joint distribution $P_{XAY}$ on $X\times A\times Y$ and let $\mathcal{E}$ denote the set of all environments. In each environment $e\in\mathcal{E}$, we have a graph dataset $G^{e}=(X^{e},A^{e},Y^{e})$, where $X^{e}\in\mathcal{X}$ are node features, $A^{e}\in\mathcal{A}$ is the adjacency matrix and $Y^{e}\in\mathcal{Y}$ is the response variable. The joint distribution $P_{XAY}^{e}$ of $X^{e}$, $A^{e}$ and $Y^{e}$ can vary across environments: $P_{XAY}^{e}\neq P_{XAY}^{e^{\prime}}$ for $e,e^{\prime}\in\mathcal{E}$ and $e\neq e^{\prime}$. In this paper, our goal is to learn a graph neural network, which can make invariant prediction across unknown environments on non-homophilous graphs, we define the node classification problem on non-homophilous graphs as:
@@
Figure 3: The module of Adaptive Neighborhood Propagation (ANP), the invariant propagation layer is proposed to combine both the high-order and low-order neighborhood information. And adaptive propagation is proposed to capture the adaptive neighborhood information, $s$ is the adaptive propagation step learned by generative network $q_{\phi}$.
@@ 3.1 Adaptive Neighborhood Propagation
In practice, there is no way to directly access the optimal propagation of neighborhood information. To capture the adaptive neighborhood information, we learn an adaptive propagation step $s$ by generative network $q_{\phi}$. Specifically, we infer the optimal propagation distribution $q_{\phi}(s_{n}|\mathbf{X},\mathbf{A})$ and learn GNN weights $\theta$ jointly with the loss function. However, the optimal propagation steps $s$ are discrete and non-differentiable which makes direct optimization difficult. Therefore, we adopt Gumbel-Softmax Sampling  Jang et al. (2016); Maddison et al. (2016), which is a simple yet effective way to substitutes the original non-differentiable sample from a discrete distribution with a differentiable sample from a corresponding Gumbel-Softmax distribution. Thus, we have the loss function as: {linenomath}
@@
where $\hat{s}$ is drawn from a categorical distribution with the discrete variational distribution $q_{{\phi}}(s_{n}|\mathbf{X},\mathbf{A})$ parameterized by $\phi$: {linenomath}
@@ 3.2 Invariant non-homophilous Graph Learning
$Assumption$: we assumes that there exists a graph representation $\Phi(X,A)$, for all environments $e,e^{\prime}\in\mathcal{E}$ , we have $P[Y^{e}|\Phi(X^{e},A^{e}))]=P[Y^{e^{\prime}}|\Phi(X^{e^{\prime}},A^{e^{\prime}}))]$.
@@
For environments $\mathcal{E}_{tr}$, each node $i$ is assigned to environment $G^{e}\in\mathcal{E}_{tr}$ with the closest cluster center by $G^{i}\mathrel{\mathop{:}}=argmin_{e}\left\|{\Psi(X_{i},A_{i})}-\mu_{e}\right\|^{2},e\in 1,2,...,K$. For every environment $G^{e}\in\mathcal{E}_{tr}$, the value of this cluster center $\mu_{e}$ is updated as $\mu_{e}\mathrel{\mathop{:}}=\frac{\sum_{i=1}^{m}l\{G^{i}=e\}{\Psi(X_{i},A_{i})}}{\sum_{i=1}^{m}l\{G^{i}=e\}}$.
@@ 4.4 Performance in Unknown Environments
For degree-related high-bias environments, We keep the nodes in the testing set unchanged and only select nodes in a specific small range of degrees as the training set, so there are strong distribution shifts between training and testing. For low-bias environments, the training and test sets are the same as shown in Appendix C. The specific statistics and detailed experimental results of low-bias and high-bias environments are shown in Appendix C. Results of low-bias and high-bias environments shown in Figure 10 show that performance of INPL and LINKX degrade under strong degree-related distribution shifts compared to the low-bias case, but the performance decline of INPL is much lower than LINKX, which proves that INPL could alleviate degree-related distribution shifts in high-bias environments. Results of more baselines and additional datasets of twitch-gamers and pokec under degree-related high-bias environments are shown in Appendix C. More results of INPL and LINKX in different degree-related bias environments are shown in Appendix C. More experiments in neighborhood pattern-related high-bias environments are in Appendix C.
Figure 10: Results of INPL and LINKX on strong distribution shifts for the task of semi-supervised node classification. INPL outperforms LINKX under strong distribution shifts and the performance improvement in the high-bias case is much higher compared to the lower-bias case.

### v2

@@ 2 Related work
Debiased Graph Neural Network. Debiased Graph Neural Networks aim to address the bias issue on graphs Chen et al. (2024); Yang et al. (2025); Lv et al. (2025); Hu et al. (2024). To overcome degree-related bias, SL-DSGCN Tang et al. (2020) mitigates the degree-related bias of GCNs by capturing both discrepancies and similarities of nodes with different degrees. ImGAGN Qu et al. (2021) is proposed to address the class-related bias, ImGAGN generates a set of synthetic minority nodes to balance the class distribution. Different from these works on a single bias, BA-GNN Chen et al. (2022) proposes a novel Bias-Aware Graph Neural Network (BA-GNN) framework by learning node representation that is invariant across different biases and distributions for invariant prediction. EERM Wu et al. (2022) facilitates graph neural networks to leverage invariance principles for prediction, EERM resorts to multiple context explorers that are adversarially trained to maximize the variance of risks from multiple virtual environments, which enables the model to extrapolate from a single observed environment which is the common case for node-level prediction. However, these works heavily rely on the assumption of homophily, which may be prohibitively failed on non-homophilous graphs. Our method differs from the above methods and aims to learn invariant graph representation for non-homophilous graphs.
@@ 3 Discovering Invariant Neighborhood Patterns
Let $G=(V,E,X)$ be a graph, where $V$ is the set of nodes, $E$ is the set of edges, $X$ $\in$ $R^{N\times d}$ is a feature matrix, each row in $X$ indicates a d-dimensional vector of a node. AA $\in$ $\{0,1\}^{N\times N}$ is the adjacency matrix, where $A_{uv}=1$ if there exists an edge between node $v_{i}$ and $v_{j}$, otherwise $A_{ij}=0$ . For semi-supervised node classification tasks, only part of nodes have known labels $Y^{o}=\{y_{1},y_{2},...y_{n}\}$, where $y_{j}\in\{0,1,...,c-1\}$ denotes the label of node $v_{j}$, $c$ is the number of classes. Similar to Chen et al. (2022), we also define a graph environment to be a joint distribution $P_{XAY}$ on $X\times A\times Y$ and let $\mathcal{E}$ denote the set of all environments. In each environment $e\in\mathcal{E}$, we have a graph dataset $G^{e}=(X^{e},A^{e},Y^{e})$, where $X^{e}\in\mathcal{X}$ are node features, $A^{e}\in\mathcal{A}$ is the adjacency matrix and $Y^{e}\in\mathcal{Y}$ is the response variable. The joint distribution $P_{XAY}^{e}$ of $X^{e}$, $A^{e}$ and $Y^{e}$ can vary across environments: $P_{XAY}^{e}\neq P_{XAY}^{e^{\prime}}$ for $e,e^{\prime}\in\mathcal{E}$ and $e\neq e^{\prime}$. In this paper, our goal is to learn a graph neural network, which can make invariant prediction across unknown environments on non-homophilous graphs, we define the node classification problem on non-homophilous graphs as:
@@
Figure 3: The module of Adaptive Neighborhood Propagation (ANP), the invariant propagation layer is proposed to combine both the high-order and low-order neighborhood information. And adaptive propagation is proposed to capture the adaptive neighborhood information, $s$ is the adaptive propagation step learned by the generative network $q_{\phi}$.
@@ 3.1 Adaptive Neighborhood Propagation
In practice, there is no way to directly access the optimal propagation of neighborhood information. To capture the adaptive neighborhood information, we learn an adaptive propagation step $s$ by generative network $q_{\phi}$. Specifically, we infer the optimal propagation distribution $q_{\phi}(s_{n}|\mathbf{X},\mathbf{A})$ and learn GNN weights $\theta$ jointly with the loss function. However, the optimal propagation steps $s$ are discrete and non-differentiable which makes direct optimization difficult. Therefore, we adopt Gumbel-Softmax Sampling  Jang et al. (2016); Maddison et al. (2016), which is a simple yet effective way to substitutes the original non-differentiable sample from a discrete distribution with a differentiable sample from a corresponding Gumbel-Softmax distribution. Thus, we have the loss function as:
@@
where $\hat{s}$ is drawn from a categorical distribution with the discrete variational distribution $q_{{\phi}}(s_{n}|\mathbf{X},\mathbf{A})$ parameterized by $\phi$:
@@ 3.2 Invariant non-homophilous Graph Learning
AssumptionAssumption: we assumes that there exists a graph representation $\Phi(X,A)$, for all environments $e,e^{\prime}\in\mathcal{E}$ , we have $P[Y^{e}|\Phi(X^{e},A^{e}))]=P[Y^{e^{\prime}}|\Phi(X^{e^{\prime}},A^{e^{\prime}}))]$.
@@
For environments $\mathcal{E}_{tr}$, each node $i$ is assigned to environment $G^{e}\in\mathcal{E}_{tr}$ with the closest cluster center by $G^{i}\mathrel{\mathop{\ordinarycolon}}=argmin_{e}\left\|{\Psi(X_{i},A_{i})}-\mu_{e}\right\|^{2},e\in 1,2,...,K$. For every environment $G^{e}\in\mathcal{E}_{tr}$, the value of this cluster center $\mu_{e}$ is updated as $\mu_{e}\mathrel{\mathop{\ordinarycolon}}=\frac{\sum_{i=1}^{m}l\{G^{i}=e\}{\Psi(X_{i},A_{i})}}{\sum_{i=1}^{m}l\{G^{i}=e\}}$.
@@ 4.4 Performance in Unknown Environments
For degree-related high-bias environments, We keep the nodes in the testing set unchanged and only select nodes in a specific small range of degrees as the training set, so there are strong distribution shifts between training and testing. For low-bias environments, the training and test sets are the same as shown in Appendix C. The specific statistics and detailed experimental results of low-bias and high-bias environments are shown in Appendix C. Results of low-bias and high-bias environments shown in Figure 10 show that the performance of INPL and LINKX degrades under strong degree-related distribution shifts compared to the low-bias case, but the performance decline of INPL is much lower than LINKX, which proves that INPL could alleviate degree-related distribution shifts in high-bias environments. Results of more baselines and additional datasets of twitch-gamers and pokec under degree-related high-bias environments are shown in Appendix C. More results of INPL and LINKX in different degree-related bias environments are shown in Appendix C. More experiments in neighborhood pattern-related high-bias environments are in Appendix C.
Figure 10: Results of INPL and LINKX on strong distribution shifts for the task of semi-supervised node classification. INPL outperforms LINKX under strong distribution shifts, and the performance improvement in the high-bias case is much higher compared to the lower-bias case.